OpenStudy (anonymous):

Find the LCM of 35z^2+175z and 7z^2+42z+35

OpenStudy (anonymous):

Start by factoring the two.

OpenStudy (anonymous):

i dont think thats right.

OpenStudy (anonymous):

Rosee, did you factor them?

OpenStudy (anonymous):

Yeah i got 35z(z+5)(z+1)

OpenStudy (anonymous):

\begin{array}{ccc}(7\cdot 5)z(z+5) & \text{and} & 7(z+5)(z+1)\end{array} Are the fully factored forms: So the LCM will be \[(7\cdot 5)z(z+1)(z+5) = 35z(z+1)(z+5)\] So yeah, that's right.